OpenStudy (anonymous):

Find the midpoint of the line segment whose endpoints are (0,1/2) and (0, 3/4).

OpenStudy (tkhunny):

\[\left(\frac{0+0}{2},\frac{½+¾}{2}\right)\] Ponder the geometry. You will see it.

OpenStudy (anonymous):

Thank You So Much ( :